### v2

@@ 5 Conclusion
This paper studies RD designs with multiple scores. We first highlight the critical but often overlooked zero-density issue with running local linear regressions using a collapsed univariate score, and then explore the use of nonparametric estimators based on forests that directly model multivariate scores. Our simulation results suggest that the zero-density issue, while hard to detect from an empirical dataset, is likely present and would impair the performance of local linear regressions. Regression forests and local linear forests, on the other hand, generally require larger sample sizes to achieve good performance at boundary points, though local linear forests are better at capturing linear trends. Although this paper does not offer a simple, universal solution to multivariate RD problems, it adds two forest-based estimators to empirical researchers’ toolkit. Overall, we recommend using the two forest-based methods as the primary approach when the sample size is large and the number of score is moderate because of their flexibility and theoretical validity. In cases where the sample size is small, local linear regressions with a collapsed univariate score can serve as a starting point, but researchers should proceed with caution and perhaps compare the results against other nonparametric alternatives introduced in this paper as a way of robustness checks.
@@ B.3 Details on the WGAN DGPs
We use the default options available from the wgan package of Athey et al. (2021) to train the WGAN for the results reported in Table 1.101010These default specifications can be found at ds-wgan.readthedocs.io/en/latest/api.html#specifications. We note that the current implementation of WGAN is only available in Python, whereas implementations of the minimax-optimal estimator and the two forests estimators are in R, and therefore we cannot, for each Monte Carlo replication, generate a new sample from WGAN in Python, keep the generated sample in memory, and then take it to R and fit the estimators: for one iteration of $1000$ Monte Carlo replications and each with a sample size of $n$, such procedure requires $n\times 1000$ observations, which quickly runs out of computer memory as we increase $n$. Instead, for each dataset that we fit a WGAN to, we draw a WGAN-generated sample of $40$ million observations, and use this large artificial sample to run the Monte Carlo simulations, treating it as a “superpopulation” from which we draw random subsamples of size $n$ used in each Monte Carlo replication. [^11] Each WGAN sample is corrected based on the geographic coordinates of the treatment boundary to ensure that for each hypothetical household, the generated coordinates match the generated treatment status. This correction is minor, affecting fewer than 2% of the observations in the WGAN sample. Figures B.2B.6 display graphical comparisons between the actual turnout, housing price and age samples and their WGAN-generated counterparts.
@@ B.5.3 Robustness to the Number of Training Epochs
In this subsection, we vary the number of training epochs specified in the WGAN algorithm for the housing price outcome as a robustness check. We consider running the algorithm for 2,000, 3,000, and 4,000 epochs, noting that 1,000 is the default option used for constructing the datasets in the main text. Each Monte Carlo trial uses a sample size of 10,000, which is moderate-to-large for RD studies. Figure B.7 shows three-dimensional surface plots that visualize the simulated CEF of housing price from the polynomial and the four WGAN DGPs, with the original observations scattered around the surfaces (i.e., CEFs). While the WGAN surfaces generally exhibit less curvature than the polynomial ones, we point out that the curvature of the treated WGAN surface (surrounded by treated dots in red) tends to increase with the number of training epochs. This instability may be attributed to the small sample size of the original data, which includes only 117 treated observations. Nevertheless, such variability in curvature could highlight certain strengths and weaknesses of the four estimators discussed in the main text. Table B.5 presents the simulation results using the three alternative DGPs. Similar to the 1,000-epoch DGP in the main text, for $n=$ 10,000, local linear regressions perform well, and the minimax-optimal estimator overcovers. [^14] Although not shown in Table B.5, we note that consistent with Panel G in Table 1, the coverage rates for local linear regressions again deteriorate as we further increase the sample size. This deterioration is possibly due to the zero-density issue of the collapsed multivariate score. Honest regression forests, however, gradually fail to capture linear signals in the treated surface near the boundary point as the surface becomes more tilted. Local linear forests alleviate this problem through the inclusion of a linear correction, as shown by their robust performance across DGPs. Nevertheless, forest-based estimators generally require larger sample sizes to achieve good coverage rates.
Figure B.7: Surface plots for simulated housing price using the polynomial DGP detailed in Section B.1.2 and the WGAN DGP introduced in Section 4.2 with 1,000, 2,000, 3,000 and 4,000 training epochs. Each point on the surfaces represents the average of 2,000 simulated housing prices at the corresponding geographic location. The outcome at the focal point is evaluated both as treated (solid red dot) and not treated (solid black dot), also averaged over 2,000 simulations. The original sample is displayed as a scatter plot around the surfaces, with red (treated) and black (control) semi-transparent dots.
@@ B.6 Hyperparameters for Regression and Local Linear Forests
For all forest-based estimators, we set the number of trees (parameter B of the regression_forest() function in the grf package) to 5,000. More trees generally reduce overfitting but at a higher computational cost. At each sufficiently large tree node, we randomly pick one feature to make the split ($\texttt{mtry}=1$). Note that in univariate designs, mtry is always 1. Small values of mtry can reduce the between-tree correlations, which would lower the overall variance. For regression forests, our choice of the subsample fraction is based on the condition in Athey et al. (2019, Theorem 5) that the subsample sizes $s_{1}$ and $s_{0}$, for the treated and control samples respectively, should satisfy $s_{1}\asymp n_{1}^{\beta}\,\,\,\text{and}\,\,\,s_{0}\asymp n_{0}^{\beta}$ for some $\beta$ lower bounded by $\beta_{min}^{RF}:=\left(1+\frac{\texttt{mtry}\times\log(1-\alpha)}{d\log(\alpha)}\right)^{-1}$, where $\alpha$ is the maximum imbalance of a split that defaults to $0.05$. Given this condition, we set the subsample fraction to $ceil[n_{j}^{\beta_{min}^{RF}}]/n_{j}$ times a constant $c\in[0.05,0.5]$ for $j\in\{0,1\}$. We leave the other tunable parameters of regression_forest() at their default values for simplicity and generalizability. [^15] See the manual for grf: regression_forest() and ll_regression_forest() for details.
Figures B.8 and B.9 show, for each DGP we consider with sample size $n=$ 5,000, how the squared bias, variance, MSE and the $95\%$ coverage rate across the 1000 Monte Carlo replications change with the scaling constant $c$ we use. As $c$ gets smaller, the bias problem and undercoverage become more pronounced, whereas the variance decreases. We recommend setting $c$ in the range of $[0.4,0.5]$ for small to moderate datasets; smaller values of $c$ can be considered with larger sample sizes. In this paper, the simulations for all forest-based estimators are conducted with the setting $c=0.4$.

### v3

@@ 6 Conclusion
This paper studies RD designs with multiple scores. We first highlight the critical but often overlooked zero-density issue with running local linear regressions using a collapsed univariate score, and then explore the use of nonparametric estimators based on forests that directly model multivariate scores. Our simulation results suggest that the zero-density issue, while hard to detect from an empirical dataset, is likely present and would impair the performance of local linear regressions. Regression forests and local linear forests, on the other hand, generally require larger sample sizes to achieve good performance at boundary points, though local linear forests are better at capturing linear trends. Although this paper does not offer a simple, universal solution to multivariate RD problems, it adds two forest-based estimators to empirical researchers’ toolkit. Overall, we recommend using the two forest-based methods as the primary approach when the sample size is large and the number of scores exceeds two because of their flexibility and theoretical validity. In cases where the sample size is small, local linear regressions with a collapsed univariate score can serve as a starting point, but researchers should proceed with caution and perhaps compare the results against other nonparametric alternatives introduced in this paper as a way of robustness checks.
@@ B.3 Details on the WGAN DGPs
We use the default options available from the wgan package of Athey et al. (2021) to train the WGAN for the results reported in Table 1.161616These default specifications can be found at wgan package manual. We note that the current implementation of WGAN is only available in Python, whereas implementations of the minimax-optimal estimator and the two forests estimators are in R, and therefore we cannot, for each Monte Carlo replication, generate a new sample from WGAN in Python, keep the generated sample in memory, and then take it to R and fit the estimators: for one iteration of $1000$ Monte Carlo replications and each with a sample size of $n$, such procedure requires $n\times 1000$ observations, which quickly runs out of computer memory as we increase $n$. Instead, for each dataset that we fit a WGAN to, we draw a WGAN-generated sample of $40$ million observations, and use this large artificial sample to run the Monte Carlo simulations, treating it as a “superpopulation” from which we draw random subsamples of size $n$ used in each Monte Carlo replication. [^17] Each WGAN sample is corrected based on the geographic coordinates of the treatment boundary to ensure that for each hypothetical household, the generated coordinates match the generated treatment status. This correction is minor, affecting fewer than 2% of the observations in the WGAN sample. Figures B.2B.6 display graphical comparisons between the actual turnout, housing price and age samples and their WGAN-generated counterparts.
@@ B.5.3 Robustness to the Number of Training Epochs
In this subsection, we vary the number of training epochs specified in the WGAN algorithm for the housing price outcome as a robustness check. We consider running the algorithm for 2,000, 3,000, and 4,000 epochs, noting that 1,000 is the default option used for constructing the datasets in the main text. Each Monte Carlo trial uses a sample size of 10,000, which is moderate-to-large for RD studies. Figure B.7 shows three-dimensional surface plots that visualize the simulated CEFs of housing price from the polynomial and the four WGAN DGPs, with the original observations scattered around the surfaces (i.e., CEFs). While the WGAN surfaces generally exhibit less curvature than the polynomial ones, we point out that the curvature of the treated WGAN surface (surrounded by treated dots in red) tends to increase with the number of training epochs. This instability may be attributed to the small sample size of the original data, which includes only 117 treated observations. Nevertheless, such variability in curvature could highlight certain strengths and weaknesses of the four estimators discussed in the main text. Table B.5 presents the simulation results using the three alternative DGPs. Similar to the 1,000-epoch DGP in the main text, for $n=$ 10,000, local linear regressions perform well, and the minimax-optimal estimator overcovers. [^20] Although not shown in Table B.5, we note that consistent with Panel G in Table 1, the coverage rates for local linear regressions again deteriorate as we further increase the sample size. This deterioration is possibly due to the zero-density issue of the collapsed multivariate score. Honest regression forests, however, gradually fail to capture linear signals in the treated surface near the boundary point as the surface becomes more tilted. Local linear forests alleviate this problem through the inclusion of a linear correction, as shown by their robust performance across DGPs. Nevertheless, forest-based estimators generally require larger sample sizes to achieve good coverage rates.
Figure B.7: Surface plots for simulated housing price using the polynomial DGP detailed in Section B.1.2 and the WGAN DGP introduced in Section 4.1.1 with 1,000, 2,000, 3,000 and 4,000 training epochs. Each point on the surfaces represents the average of 2,000 simulated housing prices at the corresponding geographic location. The outcome at the focal point is evaluated both as treated (solid red dot) and not treated (solid black dot), also averaged over 2,000 simulations. The original sample is displayed as a scatter plot around the surfaces, with red (treated) and black (control) semi-transparent dots.
@@ B.6 Hyperparameters for Regression and Local Linear Forests
For all forest-based estimators, we set the number of trees (argument B of the regression_forest() function in the grf package) to 5,000 for the simulations and 50,000 for the CARES Act empirical application. More trees generally reduce overfitting but at a higher computational cost. At each sufficiently large tree node, we randomly pick one feature to make the split ($\texttt{mtry}=1$). Note that in univariate designs, mtry is always 1. Small values of mtry can reduce the between-tree correlations, which would lower the overall variance. For regression forests, our choice of the subsample fraction is based on the condition in Athey et al. (2019, Theorem 5) that the subsample sizes $s_{1}$ and $s_{0}$, for the treated and control samples respectively, should satisfy $s_{1}\asymp n_{1}^{\beta}\,\,\,\text{and}\,\,\,s_{0}\asymp n_{0}^{\beta}$ for some $\beta$ lower bounded by $\beta_{min}^{RF}:=\left(1+\frac{\texttt{mtry}\times\log(1-\alpha)}{d\log(\alpha)}\right)^{-1}$, where $\alpha$ is the maximum imbalance of a split that defaults to $0.05$. Given this condition, we set the subsample fraction to $ceil[n_{j}^{\beta_{min}^{RF}}]/n_{j}$ times a constant $c\in[0.05,0.5]$ for $j\in\{0,1\}$. We leave the other tunable parameters of regression_forest() at their default values for simplicity and generalizability. [^21] See the grf package manual.
Figures B.8 and B.9 show, for each DGP we consider with sample size $n=$ 5,000, how the squared bias, variance, MSE and the $95\%$ coverage rate across the 1000 Monte Carlo replications change with the scaling constant $c$ we use. As $c$ gets smaller, the bias problem and undercoverage become more pronounced, whereas the variance decreases. We recommend setting $c$ in the range of $[0.4,0.5]$ for small to moderate datasets; smaller values of $c$ can be considered with larger sample sizes. In this paper, the empirical application and simulations for all forest-based estimators are conducted with $c=0.4$.